# Lusztig sheaves, characteristic cycles and the Borel-Moore homology of Nakajima’s quiver varieties

### v1

@@ Abstract.
By using characteristic cycles, we build a morphism from the canonical bases of integrable highest weight modules of quantum groups to the top Borel-Moore homology groups of Nakajima’s quiver and tensor product varieties, and compare the canonical bases and the fundamental classes. As an application, we show that Nakajima’s realization of irreducible highest weight modules and their tensor products can be defined over $\mathbb{Z}$. We also prove Nakajima’s conjecture on the canonical isomorphism of tensor product varieties.
@@ 1.1.2. Realization of integrable modules of universal enveloping algebras
Later in [24], Nakajima has considered a certain $\mathbb{C}^{\ast}$-action on $\mathfrak{M}(\omega)$ and defined his tensor product variety $\tilde{\mathfrak{Z}}$. The convolution of Hecke correspondences also gives a $\mathfrak{g}$-module structure on the top BM homology group of $\tilde{\mathfrak{Z}}$, and Nakajima has showed that $\mathbf{H}(\tilde{\mathfrak{Z}})$ is isomorphic to the tensor product $L_{1}(\lambda)\otimes L_{2}(\lambda)$. Since his proof relies on crystal structure and the calculation of characters, the isomorphism $\mathbf{H}(\tilde{\mathfrak{Z}})\cong L_{1}(\lambda)\otimes L_{2}(\lambda)$ is not canonical. Nakajima has conjectured that there is a unique canonical isomorphism satisfies certain properties. (See [24, Conjecture 5.6] or Conjecture 5.7 of the present paper for details.)
@@ 1.2. The results of the present paper
In order to apply characteristic cycle map defined by Hennecart in [9], we need to generalize the main results of [8] and [7] from varieties over $\overline{\mathbb{F}}_{q}$ to complex varieties. We consider the $\mathbf{G}_{\Omega}=\prod\limits_{h\in\Omega}\mathbb{G}^{h}_{m}$-action on the moduli spaces and define certain localizations $\mathcal{Q}_{\nu,\omega^{1}}/\mathcal{N}_{\nu}$ and $\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}/\mathcal{N}_{\nu}$ of $\mathbf{G}_{\Omega}$-equivariant Lusztig sheaves, then the Grothendieck groups $\mathcal{L}(\omega^{1})$ and $\mathcal{L}(\omega^{1},\omega^{2})$ are canonically isomorphic to irreducible highest weight modules and their tensor products. Moreover, simple perverse sheaves provide the canonical bases of these modules.
@@
As an application, we can prove Nakajima’s conjecture, which has been mentioned in Section 1.2.
@@ Acknowledgements.
Throughout this paper, all varieties are over $\mathbb{C}$ and all sheaves are of $\mathbb{C}$-vector spaces.
(1) Let $X$ be an algebraic variety. We denote by $\mathcal{D}^{b}_{c}(X)$ the bounded derived category of constructible sheaves on $X$ and denote by $\mathbf{H}^{{\rm{BM}}}_{*}(X,\mathbb{Z})$ the Borel-Moore homologies of $X$ which are normalised such that $\mathbf{H}^{{\rm{BM}}}_{2\dim X}(X,\mathbb{Z})$ has a basis given by fundamental classes of irreducible components of $X$.
@@ 2.1. Equivariant characteristic cycle and induction
In this subsection, we recall the formalism of the equivariant characteristic cycle in [9, Section 2.2] and the induction in [9, Section 4]. Let $X$ be a smooth algebraic variety and $G$ be an algebraic group acting on $X$.
@@
We refer [9, Section 2.1] for the definition of the characteristic cycle ${\rm{CC}}:\mathcal{K}_{0}(\mathcal{D}^{b}_{c,\mathcal{S}_{X}}(X))\rightarrow\mathbf{H}^{BM}_{2{\rm{dim}}X}(\Lambda,\mathbb{Z})$. We remark that we do not take the shifts $[\dim G]$ in the forgetful functor ${\rm{For}}^{G}$ which is different from [9, Section 2.2], since we use the perverse t-structure defined in [4], also see [1, Section 6.4], which is different from the perverse t-structure used in [9].
@@ 2.1.1. Equivariant induction diagram
defined by $f(m,v)=(m,p(v)),g(m,v)=mq(v)$ for any $(m,v)\in W$. Note that $f$ is smooth of relative dimension $d$, $g$ is proper and $(f,g):W\rightarrow X\times X^{\prime}$ is a closed embedding. We refer [9, Section 3.2] for details.
@@ 2.1.2. Induction of equivariant complex
We refer [9, Section 4.2] for details.
@@ 2.1.3. Induction of (equivariant) Borel-Moore homology
where $-1$ is the automorphism on $T^{\ast}X^{\prime}$ acting by $-1$ on the fibers. We refer [9, Section 3.5] for details.
@@
We refer [9, Section 4.3] for details. In particular, for a fixed $P$-invariant Whitney stratification $\mathcal{S}_{Y}$ of $Y$, we set $\Lambda_{Y}=\coprod_{S\in\mathcal{S}_{Y}}T^{\ast}_{S}Y$ and $\Lambda=G\times^{P}\Lambda_{Y}$, then there is an induction isomorphism
@@
We refer [9, Section 4.5] for details.
@@ 3.1. Lusztig sheaves
We refer [16, Section 9.1] for details of this subsection.
@@
An orientation of this graph is a subset $\Omega\subset H$ such that $\Omega\cup\overline{\Omega}=H,\Omega\cap\overline{\Omega}=\varnothing$. Given an orientation, we obtain a finite quiver $Q=(I,H,\Omega)$ without loops, where $I$ is the set of vertices, and $\Omega$ is the set of arrows such that for any $h\in\Omega$, its source and target are $h^{\prime}$ and $h^{\prime\prime}$ respectively.
@@
acts on $\mathbf{E}_{\mathbf{V},\Omega}$ by conjugation $g.x=(g_{h^{\prime\prime}}x_{h}g_{h^{\prime}}^{-1})_{h\in\Omega}$. We also define the torus
$\mathbb{T}=\mathbb{G}_{m}\times...\times\mathbb{G}_{m}$
of rank $\sharp\Omega$ such that it acts on $\mathbf{E}_{\mathbf{V},\Omega}$ by scaling $t.x=(t_{h}x_{h})_{h\in\Omega}$ when we use Fourier-Laumon transform in subsection 3.3.
Let $\mathcal{V}_{\nu}$ be the set of sequences $\boldsymbol{\nu}=(\nu_{1},\nu_{2},...,\nu_{m})$ in $\mathbb{N}I$ such that each $\nu_{l}=n_{l}i_{l}$ for some $n_{l}\in\mathbb{N},i_{l}\in I$ and $\sum_{l=1}^{m}\nu_{l}=\nu$. For any $\boldsymbol{\nu}=(\nu_{1},\nu_{2},...,\nu_{m})\in\mathcal{V}_{\nu}$, let ~$\tilde{\mathcal{F}}_{\boldsymbol{\nu}}$ be the variety of pairs $(x,f)$, where $x\in\mathbf{E}_{\mathbf{V},\Omega}$ and $f=(\mathbf{V}=\mathbf{V}^{0}\supseteq\mathbf{V}^{1}\supseteq...\supseteq\mathbf{V}^{m}=0)$ is a sequence of $I$-graded subspaces of $\mathbf{V}$ satisfying $|\mathbf{V}^{l-1}/\mathbf{V}^{l}|=\nu^{l}$ and $x_{h}(\mathbf{V}^{l}_{h^{\prime}})\subseteq\mathbf{V}^{l}_{h^{\prime\prime}}$ for any $h\in\Omega,l=1,...,m$, then $\mathbf{G}_{\mathbf{V}}$ acts on ~$\tilde{\mathcal{F}}_{\boldsymbol{\nu}}$ by $g.(x,f)=(g.x,g.f)$, where $g.f=(\mathbf{V}=g\mathbf{V}^{0}\supseteq g\mathbf{V}^{1}\supseteq...\supseteq g\mathbf{V}^{m}=0)$, and $\mathbb{T}$ acts on ~$\tilde{\mathcal{F}}_{\boldsymbol{\nu}}$ by $t.(x,f)=(t.x,f)$.
Lemma 3.1 ([14, Lemma 1.6]).
@@
(b) The natural projection $\pi_{\boldsymbol{\nu}}:\tilde{\mathcal{F}}_{\boldsymbol{\nu}}\rightarrow\mathbf{E}_{\mathbf{V},\Omega}$ is proper and $\mathbf{G}_{\mathbf{V}}\times\mathbb{T}$-equivariant.
By the decomposition theorem, see [4, Theorem 5.3], the complex
@@
is a $\mathbf{G}_{\mathbf{V}}\times\mathbb{T}$-equivariant semisimple complex on $\mathbf{E}_{\mathbf{V},\Omega}$. Let $\mathcal{P}_{\nu}$ be the full subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V},\Omega})$ consisting of simple perverse sheaves $L$ such that $L[n]$ is a direct summand of $L_{\boldsymbol{\nu}}$ for some $n\in\mathbb{Z},\boldsymbol{\nu}\in\mathcal{S}_{\nu}$, and let $\mathcal{Q}_{\nu}$ be the full subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V},\Omega})$ consisting of direct sums of complexes of the form $L[n]$ for any $L\in\mathcal{P}_{\nu},n\in\mathbb{Z}$. We call $\mathcal{Q}_{\nu}$ the category of Lusztig sheaves.
@@ 3.2. Induction and restriction functors
We refer [16, Section 9.2] for details of this subsection.
For any $\nu=\nu^{\prime}+\nu^{\prime\prime}\in\mathbb{N}I$ and $I$-graded $\mathbb{C}$-vector spaces $\mathbf{V}=\mathbf{V}^{\prime}\oplus\mathbf{V}^{\prime\prime}$ such that $|\mathbf{V}|=\nu,|\mathbf{V}^{\prime}|=\nu^{\prime},|\mathbf{V}^{\prime\prime}|=\nu^{\prime\prime}$, let $F$ be the closed subset of $\mathbf{E}_{\mathbf{V},\Omega}$ consisting of $x$ such that $x_{h}(V_{h^{\prime}}^{\prime\prime})\subseteq V_{h^{\prime\prime}}^{\prime\prime}$ for any $h\in\Omega$. For any $x\in F$, there are natural linear maps $x^{\prime}_{h}:\mathbf{V}^{\prime}_{h^{\prime}}\rightarrow\mathbf{V}^{\prime}_{h^{\prime\prime}},x^{\prime\prime}_{h}:\mathbf{V}^{\prime\prime}_{h^{\prime}}\rightarrow\mathbf{V}^{\prime\prime}_{h^{\prime\prime}}$ induced by $x_{h}:\mathbf{V}_{h^{\prime}}\rightarrow\mathbf{V}_{h^{\prime\prime}}$ for any $h\in\Omega$, and then we obtain elements $x^{\prime}=(x^{\prime}_{h})_{h\in\Omega}\in\mathbf{E}_{\mathbf{V}^{\prime},\Omega},x^{\prime\prime}=(x^{\prime\prime}_{h})_{h\in\Omega}\in\mathbf{E}_{\mathbf{V}^{\prime\prime},\Omega}$. Consider the following diagram
@@
where $p_{1}(g,x)=\kappa(x),p_{2}(g,x)=(g,x),p_{3}(g,x)=\iota(x)$.
@@
$\mathbf{Ind}^{\mathbf{V}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}}:$ $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{\prime}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{\prime},\Omega})\boxtimes\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{\prime\prime}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{\prime\prime},\Omega})\rightarrow\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V},\Omega}),$
@@
$\mathbf{Res}^{\mathbf{V}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}}:$ $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V},\Omega})\rightarrow\mathcal{D}^{b}_{G_{\mathbf{V}^{\prime}}\times\mathbb{T}\times G_{\mathbf{V}^{\prime\prime}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}^{\prime},\Omega}\times\mathbf{E}_{\mathbf{V}^{\prime\prime},\Omega}),$
$C\mapsto(\kappa_{\Omega})_{!}(\iota_{\Omega})^{\ast}(C)[-\langle\nu^{\prime},\nu^{\prime\prime}\rangle_{Q}],$

### v3

@@ Abstract.
By using characteristic cycles, we build a morphism from the canonical bases of integrable highest weight modules of quantum groups to the top Borel-Moore homology groups of Nakajima’s quiver and tensor product varieties, and compare the canonical bases and the fundamental classes. As an application, we show that Nakajima’s realization of irreducible highest weight modules and their tensor products can be defined over $\mathbb{Z}$. We also give a new proof of Nakajima’s conjecture on the canonical isomorphism of tensor product varieties.
@@ 1.1.2. Realization of integrable modules of universal enveloping algebras
Later in [28], Nakajima has considered a certain $\mathbb{C}^{\ast}$-action on $\mathfrak{M}(\omega)$ and defined his tensor product variety $\tilde{\mathfrak{Z}}$. The convolution of Hecke correspondences also gives a $\mathfrak{g}$-module structure on the top BM homology group of $\tilde{\mathfrak{Z}}$, and Nakajima has showed that $\mathbf{H}(\tilde{\mathfrak{Z}})$ is isomorphic to the tensor product $L_{1}(\lambda)\otimes L_{2}(\lambda)$. Since his proof relies on crystal structure and the calculation of characters, the isomorphism $\mathbf{H}(\tilde{\mathfrak{Z}})\cong L_{1}(\lambda)\otimes L_{2}(\lambda)$ is not canonical. Nakajima has conjectured that there is a unique canonical isomorphism satisfies certain properties. (See [28, Conjecture 5.6] or Conjecture 5.7 of the present paper for details.) This conjectural canonical isomorphism has been realized by the stable envelope in [24].
@@ 1.2. The results of the present paper
In order to apply characteristic cycle map defined by Hennecart in [14], we need to generalize the main results of [12] and [10] from varieties over $\overline{\mathbb{F}}_{q}$ to complex varieties. We consider the $\mathbb{T}=\prod\limits_{i\in I}\mathbb{G}^{i}_{m}$-action on the moduli spaces and define certain localizations $\mathcal{Q}_{\nu,\omega^{1}}/\mathcal{N}_{\nu}$ and $\mathcal{Q}_{\nu,\omega^{1},\omega^{2}}/\mathcal{N}_{\nu}$ of $\mathbf{G}_{\mathbf{V}}\times\mathbb{T}$-equivariant Lusztig sheaves, then the Grothendieck groups $\mathcal{L}(\omega^{1})$ and $\mathcal{L}(\omega^{1},\omega^{2})$ are canonically isomorphic to irreducible highest weight modules and their tensor products. Moreover, simple perverse sheaves provide the canonical bases of these modules.
@@
As an application, we can give a new proof of Nakajima’s conjecture, which has been mentioned in Section 1.2.
@@ Acknowledgements.
Throughout this paper, all smooth varieties are over $\mathbb{C}$ with complex analytic topology and all sheaves are of $\mathbb{C}$-vector spaces.
(1) Let $X$ be a smooth variety. We denote by $\mathcal{D}^{b}_{c}(X)$ the bounded derived category of constructible sheaves on $X$ and denote by $\mathbf{H}^{{\rm{BM}}}_{*}(X,\mathbb{Z})$ the Borel-Moore homologies of $X$ which are normalised such that $\mathbf{H}^{{\rm{BM}}}_{2\dim X}(X,\mathbb{Z})$ has a basis given by fundamental classes of irreducible components of $X$.
@@ 2.1. Equivariant characteristic cycle and induction
In this subsection, we recall the formalism of the equivariant characteristic cycle in [14, Section 2.2] and the induction in [14, Section 4]. Let $X$ be a smooth algebraic variety and $G$ be an algebraic group acting on $X$.
@@
We refer [14, Section 2.1] for the definition of the characteristic cycle ${\rm{CC}}:\mathcal{K}_{0}(\mathcal{D}^{b}_{c,\mathcal{S}_{X}}(X))\rightarrow\mathbf{H}^{BM}_{2{\rm{dim}}X}(\Lambda,\mathbb{Z})$. We remark that we do not take the shifts $[\dim G]$ in the forgetful functor ${\rm{For}}^{G}$ which is different from [14, Section 2.2], since we use the perverse t-structure defined in [5], also see [1, Section 6.4], which is different from the perverse t-structure used in [14].
@@ 2.1.1. Equivariant induction diagram
defined by $f(m,v)=(m,p(v)),g(m,v)=mq(v)$ for any $(m,v)\in W$. Note that $f$ is smooth of relative dimension $d$, $g$ is proper and $(f,g):W\rightarrow X\times X^{\prime}$ is a closed embedding. We refer [14, Section 3.2] for details.
@@ 2.1.2. Induction of equivariant complex
We refer [14, Section 4.2] for details.
@@ 2.1.3. Induction of (equivariant) Borel-Moore homology
where $-1$ is the automorphism on $T^{\ast}X^{\prime}$ acting by $-1$ on the fibers. We refer [14, Section 3.5] for details.
@@
We refer [14, Section 4.3] for details. In particular, for a fixed $P$-invariant Whitney stratification $\mathcal{S}_{Y}$ of $Y$, we set $\Lambda_{Y}=\coprod_{S\in\mathcal{S}_{Y}}T^{\ast}_{S}Y$ and $\Lambda=G\times^{P}\Lambda_{Y}$, then there is an induction isomorphism
@@
We refer [14, Section 4.5] for details.
@@ 3.1. Lusztig sheaves
We refer [20, Section 9.1] for details of this subsection.
@@
An orientation of this graph is a subset $\Omega\subset H$ such that $\Omega\cup\overline{\Omega}=H,\Omega\cap\overline{\Omega}=\varnothing$. Given an orientation, we obtain a finite quiver $Q=(I,H,\Omega)$ without loops, where $I$ is the set of vertices, and $\Omega$ is the set of arrows such that for any $h\in\Omega$, its source and target are $h^{\prime}$ and $h^{\prime\prime}$ respectively. We assume $Q$ doesn’t have loops and oriented cycles, but may have multiple arrows.
@@
acts on $\mathbf{E}_{\mathbf{V},\Omega}$ by conjugation $g.x=(g_{h^{\prime\prime}}x_{h}g_{h^{\prime}}^{-1})_{h\in\Omega}$.
Let $\mathcal{V}_{\nu}$ be the set of sequences $\boldsymbol{\nu}=(\nu_{1},\nu_{2},...,\nu_{m})$ in $\mathbb{N}I$ such that each $\nu_{l}=n_{l}i_{l}$ for some $n_{l}\in\mathbb{N},i_{l}\in I$ and $\sum_{l=1}^{m}\nu_{l}=\nu$. For any $\boldsymbol{\nu}=(\nu_{1},\nu_{2},...,\nu_{m})\in\mathcal{V}_{\nu}$, let ~$\tilde{\mathcal{F}}_{\boldsymbol{\nu}}$ be the variety of pairs $(x,f)$, where $x\in\mathbf{E}_{\mathbf{V},\Omega}$ and $f=(\mathbf{V}=\mathbf{V}^{0}\supseteq\mathbf{V}^{1}\supseteq...\supseteq\mathbf{V}^{m}=0)$ is a sequence of $I$-graded subspaces of $\mathbf{V}$ satisfying $|\mathbf{V}^{l-1}/\mathbf{V}^{l}|=\nu^{l}$ and $x_{h}(\mathbf{V}^{l}_{h^{\prime}})\subseteq\mathbf{V}^{l}_{h^{\prime\prime}}$ for any $h\in\Omega,l=1,...,m$, then $\mathbf{G}_{\mathbf{V}}$ acts on ~$\tilde{\mathcal{F}}_{\boldsymbol{\nu}}$ by $g.(x,f)=(g.x,g.f)$, where $g.f=(\mathbf{V}=g\mathbf{V}^{0}\supseteq g\mathbf{V}^{1}\supseteq...\supseteq g\mathbf{V}^{m}=0)$.
Lemma 3.1 ([18, Lemma 1.6]).
@@
(b) The natural projection $\pi_{\boldsymbol{\nu}}:\tilde{\mathcal{F}}_{\boldsymbol{\nu}}\rightarrow\mathbf{E}_{\mathbf{V},\Omega}$ is proper and $\mathbf{G}_{\mathbf{V}}$-equivariant.
By the decomposition theorem, see [5, Theorem 5.3], the complex
@@
is a $\mathbf{G}_{\mathbf{V}}$-equivariant semisimple complex on $\mathbf{E}_{\mathbf{V},\Omega}$. Let $\mathcal{P}_{\nu}$ be the full subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega})$ consisting of simple perverse sheaves $L$ such that $L[n]$ is a direct summand of $L_{\boldsymbol{\nu}}$ for some $n\in\mathbb{Z},\boldsymbol{\nu}\in\mathcal{S}_{\nu}$, and let $\mathcal{Q}_{\nu}$ be the full subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega})$ consisting of direct sums of complexes of the form $L[n]$ for any $L\in\mathcal{P}_{\nu},n\in\mathbb{Z}$. We call $\mathcal{Q}_{\nu}$ the category of Lusztig sheaves.
@@ 3.2. Induction and restriction functors
We refer [20, Section 9.2] for details of this subsection.
For any $\nu=\nu^{\prime}+\nu^{\prime\prime}\in\mathbb{N}I$ and $I$-graded $\mathbb{C}$-vector spaces $\mathbf{V}=\mathbf{V}^{\prime}\oplus\mathbf{V}^{\prime\prime}$ such that $|\mathbf{V}|=\nu,|\mathbf{V}^{\prime}|=\nu^{\prime},|\mathbf{V}^{\prime\prime}|=\nu^{\prime\prime}$, let $F$ be the closed subset of $\mathbf{E}_{\mathbf{V},\Omega}$ consisting of $x$ such that $x_{h}(\mathbf{V}_{h^{\prime}}^{\prime\prime})\subseteq\mathbf{V}_{h^{\prime\prime}}^{\prime\prime}$ for any $h\in\Omega$. For any $x\in F$, there are natural linear maps $x^{\prime}_{h}:\mathbf{V}^{\prime}_{h^{\prime}}\rightarrow\mathbf{V}^{\prime}_{h^{\prime\prime}},x^{\prime\prime}_{h}:\mathbf{V}^{\prime\prime}_{h^{\prime}}\rightarrow\mathbf{V}^{\prime\prime}_{h^{\prime\prime}}$ induced by $x_{h}:\mathbf{V}_{h^{\prime}}\rightarrow\mathbf{V}_{h^{\prime\prime}}$ for any $h\in\Omega$, and then we obtain elements $x^{\prime}=(x^{\prime}_{h})_{h\in\Omega}\in\mathbf{E}_{\mathbf{V}^{\prime},\Omega},x^{\prime\prime}=(x^{\prime\prime}_{h})_{h\in\Omega}\in\mathbf{E}_{\mathbf{V}^{\prime\prime},\Omega}$. Consider the following diagram
@@
where $p_{1}(g,x)=\kappa(x),p_{2}(g,x)=(g,x),p_{3}(g,x)=g_{\cdot}\iota(x)$.
@@
$\mathbf{Ind}^{\mathbf{V}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}}:$ $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{\prime}}}(\mathbf{E}_{\mathbf{V}^{\prime},\Omega})\boxtimes\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}^{\prime\prime}}}(\mathbf{E}_{\mathbf{V}^{\prime\prime},\Omega})\rightarrow\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega}),$
@@
$\mathbf{Res}^{\mathbf{V}}_{\mathbf{V}^{\prime},\mathbf{V}^{\prime\prime}}:$ $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V},\Omega})\rightarrow\mathcal{D}^{b}_{G_{\mathbf{V}^{\prime}}\times G_{\mathbf{V}^{\prime\prime}}}(\mathbf{E}_{\mathbf{V}^{\prime},\Omega}\times\mathbf{E}_{\mathbf{V}^{\prime\prime},\Omega}),$
$C\mapsto(\kappa)_{!}(\iota)^{\ast}(C)[-\langle\nu^{\prime},\nu^{\prime\prime}\rangle_{Q}],$